1. Let p-: P(head) when a coin is tossed, and consider the hypotheses Ho : p-0.5 and H. : p-0.1. Suppose the coin is tossed 10 times, and te critcal region is defined by get 2 heads or less. That is the null hypothesis is rejected if the number of observed heads is 2 or less. (a) Find the significance level of the test (a). (b) Find the power of the test (1 -B).
Answer #1

a)P(significance level) =P(getting 2 or less heads given p=0.5)=P(X<=2)=P(X=0)+P(X=1)+P(X=2)

=10C0(0.5)0(0.5)10+10C1(0.5)1(0.5)9+10C2(0.5)2(0.5)8=0.0547

b)

P(power of the test ) =P(getting 2 or less heads given p=0.1)=P(X<=2)=P(X=0)+P(X=1)+P(X=2)

=10C0(0.1)0(0.9)10+10C1(0.1)1(0.9)9+10C2(0.1)2(0.9)8=0.9298